OpenStudy (anonymous):

what number must you add to complete the square? x^2+8x=-3

OpenStudy (anonymous):

complete the square ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I don't know...

OpenStudy (anonymous):

u need to add 16 on both sides to complete the square (x+4)^2=13

OpenStudy (anonymous):

so just 16 no squareroot?

OpenStudy (anonymous):

wy do you need square root...? see basically the logic is to bring it to the form (x+b)^2=x^2+b2+2xb, so for getting this factorisation, see appropriately what to add. observe that the x term has co-efficient 2b and so if u hav a b^2 term u will form the complete square by the formula rite? so u add b^2=(8/2)^2=4^2=16.